Find two consecutive even integers whose sum is 50.
Thank you for your help.

Answer by checkley77(12844)

Let x & x+2 be the 2 even integers.
x+x+2=50
2x=50-2
2x=48
x=48/2
x=24 ans. for the smaller integer.
24+2=26 for the larger integer.
Proof:
24+26=50
50=50